OpenStudy (anonymous):

Solve the following systems by any method x + 3y = 12 x = y - 8

OpenStudy (aihberkhan):

Okay so since we know x is \(y - 8\) we will plug that in for x in the first equation.

OpenStudy (aihberkhan):

So you will get \(y - 8 + 3y = 12\)

OpenStudy (aihberkhan):

Now when solving you should get \(y = 5\)

OpenStudy (aihberkhan):

Now plug that 5 in for y in the second equation. So you should get \(x = 5 - 8\) So \(x = -3\)

OpenStudy (aihberkhan):

So your answer is \(x = -3\) \(y = 5\)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @sarahjoy28